It takes a girl 48 minutes to deliver newspapers on her route; however, if her brother helps, it takes them only 12 minutes. How long would it take her brother to deliver the newspapers by himself?
Answer by nerdybill(7384)

It takes a girl 48 minutes to deliver newspapers on her route; however, if her brother helps, it takes them only 12 minutes. How long would it take her brother to deliver the newspapers by himself?
.
Let x = time it takes the brother to deliver by himself
then
12(1/48 + 1/x) = 1
Multiply both sides by 48x:
12(x + 48) = 48x
x + 48 = 4x
48 = 3x
16 minutes = x